OpenStudy (anonymous):

–3x + 2y = 8 change the equation into slope-intercept form

OpenStudy (anonymous):

First you add -3x to 8. That will give you 2y=3x+8. Then you divide by 2 on all three of those. What did you get?

OpenStudy (anonymous):

-3x + 2y = 8 -3x + 3x + 2y = 3x + 8 2y/2= 3x/2 + 8/2 Y = x + 4 Is this correct? I dont think it is!

OpenStudy (anonymous):

y=4+3/2

OpenStudy (anonymous):

y=3/2x+4